OpenStudy (eric_d):

Use the expansion of (2-x)^5 to evaluate (1.98)^5, correct to 5 d.p

OpenStudy (anonymous):

1.98 = (2-0.2)

OpenStudy (eric_d):

ok..

OpenStudy (anonymous):

so u can solve it now?

OpenStudy (eric_d):

What do I need to do

OpenStudy (anonymous):

You have a formula for (x-y)^5 ?

OpenStudy (anonymous):

u can also write it as (2-0.02)^5 which is (2-0.02)^2 * (2-0.02)^2 * (2-0.02)

OpenStudy (anonymous):

right?

OpenStudy (eric_d):

why ^2

OpenStudy (anonymous):

Total power is five. You divide the powers into 2, 2 and 1

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

@Brainybeauty

OpenStudy (eric_d):

@OOOPS

OpenStudy (anonymous):

do you know Pascal's triangle?

OpenStudy (eric_d):

That's a longer method

OpenStudy (anonymous):

ok, so you want taking Brainy beauty 's one as above?? just apply

OpenStudy (eric_d):

Yes.. but how

OpenStudy (anonymous):

you have 3 terms, right? first 2 terms are the same (2-0.02)^2 hehehe it is (a-b)^2 = a^2 -2ab +b^2 , your a =2, your b = 0.02, just apply to get (2-0.02)^2

OpenStudy (anonymous):

\[(\color{red}{a}-\color{blue}{b})^2=\color{red}{a^2}-2*\color{red}{a}*\color{blue}{b}+\color{blue}{b^2}\\(\color{red}{2}-\color{blue}{0.02})^2=\color{red}{2^2}-2*\color{red}{2}*\color{blue}{0.02}+\color{blue}{0.02^2}\]

OpenStudy (anonymous):

you do the rest, please.

OpenStudy (eric_d):

Okay! Got it!

OpenStudy (eric_d):

Thank you @Brainybeauty and @OOOPS